OpenStudy (ohohaye):

Please help me What is the period for the secant function? (Answer in radians.)

OpenStudy (anonymous):

Do you know what the period for the cosecant function is? @ohohaye

OpenStudy (ohohaye):

Not really

OpenStudy (anonymous):

Do you know what the period for the sin function is then?

OpenStudy (ohohaye):

No, I feel like I do, but I forgot what it was

OpenStudy (anonymous):

Well, I'll tell you.

OpenStudy (anonymous):

So the period of a \(\sf sine~ function ~is ~2\pi\) \(\sf cosine~ function ~is ~2\pi\) \(\sf tangent~ function ~is ~\pi\)

OpenStudy (ohohaye):

oh ok, yeah that seems familiar to me

OpenStudy (anonymous):

And so the periods will be: \(\sf cosecant~ function ~is ~2\pi\) \(\sf secant~ function ~is ~2\pi\) \(\sf cotangent~ function ~is ~2\pi\)

OpenStudy (ohohaye):

ok, thank you

OpenStudy (anonymous):

Anytime :)